OpenStudy (evansanchez):

The graph of which of the following equations contains the points (2, 3) and (3, 2)? y = x - 1 y = 5 - x y= 7 - 2x

OpenStudy (eliesaab):

Replace and decide

OpenStudy (evansanchez):

not good at this so idk.

OpenStudy (eliesaab):

take y= x-1, if x=2, how much is y?

OpenStudy (evansanchez):

y=1 ? @eliesaab

OpenStudy (eliesaab):

So the first choice is out since (2,3) is not on the line

OpenStudy (eliesaab):

Let us try y=5- x to see if the points (2, 3) and (3, 2)

OpenStudy (evansanchez):

what would x be?

OpenStudy (eliesaab):

for y=5-x, if x =2, how much is y?

OpenStudy (evansanchez):

y=3

OpenStudy (eliesaab):

So the point (2,3) is on the line 5-x

OpenStudy (eliesaab):

Let us see if the second point (3,2) is on it if y=5-x and x=3, how much is y?

OpenStudy (evansanchez):

y=2

OpenStudy (eliesaab):

So the second point (3,2) is on y=5-x. So what is the answer to your problem?

OpenStudy (evansanchez):

B.

OpenStudy (eliesaab):

Yes

OpenStudy (evansanchez):

thank you!!

OpenStudy (eliesaab):

Did you understand it?

OpenStudy (eliesaab):

YW

OpenStudy (evansanchez):

yes, you helped a lot

OpenStudy (eliesaab):

Great